OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 78.54 km2. A. 5 ft B. 15 ft C. 10 ft D. 2.5 ft

hartnn (hartnn):

surface area of sphere is given by \(\large S=4\pi r^2\) here, S =78.54 can you find r ??

OpenStudy (anonymous):

how do find R?

OpenStudy (anonymous):

i*

hartnn (hartnn):

by dividing both sides by 4pi, first and then taking square root.

OpenStudy (anonymous):

I don't think i'm doing this right -_- i got like 6.25 and so on

hartnn (hartnn):

you gor r^2 = 6.25 that is correct!

hartnn (hartnn):

now simply take square root

OpenStudy (anonymous):

i don't have my calculator so i cant do it in my head -_- lmao

hartnn (hartnn):

ok, \(if, r^2=6.25,r=2.5\)

OpenStudy (anonymous):

nvm googled a calculator lol

OpenStudy (anonymous):

thank you, i got it now ;b

hartnn (hartnn):

haha, even i did it with google :P